40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10562     Entry time: Fri Oct 3 03:02:17 2014
Author: ericq 
Type: Update 
Category: LSC 
Subject: No luck locking DRMI 

I haven't been able to lock the DRMI tonight, neither with 1F and no arms nor 3F and arms held off with ALS... I tried previous recipes, and new combinations informed by simulations I've run, to no avail. 

I touched the alignment of the green beat PD on the PSL table, since the X beatnote was rather low, but wasn't able to improve it by much. I never took a spectrum, since it wasn't my main focus tonight, but the low frequency motion of both arms on ALS, as observed by RIN, was good as I've ever seen it. 

In our WFS work earlier today, Koji and I reset the WFS offsets, and it actually seems to have helped a good deal, in terms of the "fuzz" of MC REFL on the wall striptool. I had previously presumed this to be due to excess angular motion, but perhaps it is more accurately described as an alignment offset that let the nominal angular motion couple into the RIN more. 

ELOG V3.1.3-